how do i decide whether the graphs of the two equations are parallel lines
y=2x-12
6x+3y=0

Answer by filboi(35)

The two equations are parallel lines if their slopes are equal. Meaning: if they have the same slope, they do not intersect at any point and therefore parallel to each other.
First, change the equations to the format y=mx + b. In this equation, m is the slope.
y=2x-12; m=2
6x+3y=0 ==> 3y=-6x ==> y=-2x; m=-2
You will notice that the values of the slope are equal but different in sign. Therefore, they are two intersecting lines and are in fact perpendicular to each other.